Sir please help me to factorise: (x^3+8/x) If anyone now answer please tell it

Grade:9

1 Answers

DASARI NAGESWARA RAO
6 years ago
x^3+8/x=x^3+2^3/(3rd root x)^3=x^3+(2/3rd root x)^3a^3+b^3=(a+b)(a^2-ab+b^2)=(x+2/3rd root x)(x^2+-x.2/3rd root x+(2/3rx root x)^2=(x+2/3rd root x)(x^2-2.x^2/3+4/x^2/3)